### v1

# Categorification of generic Su-Zhang Character Formula

Categorification of generic Su-Zhang Character Formula
Shunsuke Hirota
(December 18, 2025)
For semisimple Lie algebras, the BGG resolution is commonly interpreted as a categorification of the Weyl character formula. For general linear Lie superalgebras, the Kac–Wakimoto character formula is regarded as a good character formula for a certain class of ”singular” weights (the so-called Kac–Wakimoto weights). However, the combinatorial factor $\frac{1}{\operatorname{atyp}(\lambda)!}$ makes a direct BGG-type categorification of this formula rather elusive.
In this paper we construct, for a certain class of “generic” weights, a resolution that categorifies a known Weyl-type finite-sum character formula of the same flavour as the Kac–Wakimoto formula. Our resolution is built from taking images of canonical homomorphisms between Verma modules corresponding to non-conjugate Borel subalgebras via odd reflections.
Note that, while the infinite resolution of Kac–Wakimoto simple modules by Kac modules constructed by Brundan–Stroppel [brundan2010highestII] does not directly generalize the classical BGG resolution, the construction developed here does.
1 Introduction
General linear Lie superalgebras $\mathfrak{gl}(m|n)$ are widely regarded as natural generalizations of general linear Lie algebras, and their representation theory has been extensively studied as a rich and interesting subject. One of the most fundamental problems is to determine the characters of finite-dimensional simple representations. This problem was solved independently by Serganova [serganova1996kazhdan] and Brundan [brundan2003kazhdan]. Their results can be expressed as the following infinite sum:
$\operatorname{ch}L^{()}(\lambda)=\sum_{\mu\in P^{+}}K_{\lambda,\mu}(-1)\,\operatorname{ch}K_{\geq 0}(\mu),$
where $L^{()}(\lambda)$ denotes the simple highest weight module, $K_{\lambda,\mu}(-1)$ is the value at 1-1 of Serganova-Brundan’s Kazhdan–Lusztig-type polynomial (now well understood in terms of the arc–diagram calculus of Brundan–Stroppel [stroppel2012highest]), and $K_{\geq 0}(\mu)$ is the Kac module, i.e. the standard object in the finite-dimensional weight module category $\mathcal{F}$.
Building on Brundan’s result [brundan2003kazhdan], Su and Zhang [su2007character] obtained, for the first time, a Weyl-type character formula given by a finite sum. Unlike in the purely even Lie algebra case, however, the Su–Zhang character formula is rather intricate and not very convenient for explicit computations. We refer to [gorelik2023gruson] for a detailed overview of the history of character formulas for Lie superalgebras. Nevertheless, in several important classes one can obtain very neat closed character formulas [chmutov2014weyl]. The best-known example is the Kac–Wakimoto character formula (conjectured in [kac1994integrable] and proved in [chmutov2015kac, cheng2017kac, gorelik2015characters]). Very roughly, for a so-called Kac–Wakimoto weight (a.k.a. Kostant weight or totally disconnected weight) $\lambda$ one has
$\operatorname{ch}L^{()}(\lambda)=\frac{1}{\operatorname{atyp}(\lambda)!}\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})},$
where $M^{()}(w\cdot\lambda)$ is the Verma module, $\Gamma_{\lambda}$ denotes the set of atypical odd positive roots for $\lambda$ and $\operatorname{atyp}(\lambda)$ denotes the atypicality of $\lambda$.
In the classical setting of semisimple Lie algebras, the Weyl character formula is categorified by the BGG resolution [bgg1973schubert]. It is thus natural to ask whether one can construct a resolution that categorifies the Kac–Wakimoto character formula. The presence of the factor $\frac{1}{\operatorname{atyp}(\lambda)!}$, however, makes such a categorification problem rather delicate. Note that Brundan–Stroppel [brundan2010highestII] constructed, generalizing [cheng2008bgg], for Kac–Wakimoto weights $\lambda$, an explicit and very interesting resolution of the Kac module $K_{\geq 0}(\lambda)$ in terms of diagram algebras. This should be viewed as a categorification of the Serganova–Brundan infinite-sum formula, rather than of a Weyl-type formula. See also [coulembier2014bernstein].
Kac–Wakimoto weights can be considered, in a sense, as the most singular among regular dominant weights. On the other hand, there are natural “generic” classes of weights. In this paper we work with the class of $\mathfrak{g}_{-1}$-generic weights. Our notion of $\mathfrak{g}_{-1}$-genericity coincides with the $\Gamma^{+}$–genericity of Coulembier [coulembier2016bott], and is therefore weaker than the notion of weakly generic and generic in Penkov-Serganova [coulembier2016primitive, penkov1989cohomology, penkov1994generic]. In particular, any weight lying outside a small neighbourhood of the walls of the Weyl chambers is $\mathfrak{g}_{-1}$–generic.
Su and Zhang [su2007character] showed that, on a certain generic region which contains all $\mathfrak{g}_{-1}$-generic weights, their character formula simplifies to
$\operatorname{ch}L^{()}(\lambda)=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
Here the troublesome factor $\frac{1}{\operatorname{atyp}(\lambda)!}$ disappears. (See also [chmutov2014weyl] for a nice exposition.) This makes it natural to try to construct a BGG-type resolution for such $\lambda$. More precisely, we aim to construct a “narrow” Verma module $N^{()}(\lambda)$ whose character is given by
$\operatorname{ch}N^{()}(\lambda)=\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})},$
and then to build a resolution of $L^{()}(\lambda)$ in terms of the modules.
Our main point of view is that the above narrow Verma modules and their resolutions can be constructed from odd reflections and changes of Borel subalgebras. In contrast to the situation for semisimple Lie algebras, a basic classical Lie superalgebra admits several non-conjugate Borel subalgebras, and passing from one Borel to another via odd reflections is a genuinely new and essential feature of the theory. From the perspective of Nichols algebras of diagonal type [heckenberger2009classification, andruskiewitsch2017finite] and their classification via their Weyl groupoids (a.k.a. generalized root systems) [heckenberger2008generalization, heckenberger2020hopf, bonfert2024weyl], these changes of Borel are also of independent interest.
Our narrow Verma modules are defined as follows. Let $()$ be the distinguished Borel subalgebra and $(n^{m})$ the antidistinguished Borel subalgebra of $\mathfrak{gl}(m|n)$. We define the narrow Verma module $N^{()}(\lambda)$ to be the image of up to scalar, a unique nonzero homomorphism Verma modules sharing same character with different choice of Borels;
$N^{()}(\lambda):=\operatorname{Im}\!\Bigl(M^{()}(\lambda)\longrightarrow M^{(n^{m})}\bigl(\lambda-2\rho_{\overline{1}}\bigr)\Bigr).$
Our main result can be stated as follows.
Theorem 1.1.
Let $\mathfrak{g}=\mathfrak{gl}(m|n)$ and assume that $\lambda$ is a dominant integral and $\mathfrak{g}_{-1}$-generic weight. Then there exists an exact sequence of $\mathfrak{g}$-modules
$0\to N^{()}(w_{0}\cdot\lambda)\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=\ell(w_{0})-1\end{subarray}}N^{()}(w\cdot\lambda)\to\cdots\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=1\end{subarray}}N^{()}(w\cdot\lambda)\to N^{()}(\lambda)\to L^{()}(\lambda)\to 0.$
Moreover,
$\operatorname{ch}N^{()}(\lambda)=\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
In particular, if $\lambda$ is antidominant, then $N^{()}(\lambda)$ is simple. Hence we obtain a closed character formula for all $\mathfrak{g}_{-1}$-generic antidominant simple highest weight modules.
Our main tools in this paper is the restriction functor and the (dual) Kac functor. We derive the character formula for $N^{()}(\lambda)$ from the Su–Zhang character formula.
1.1 Acknowledgments
The author would like to thank Yoshiyuki Koga for suggesting this project, for computing several low-rank examples, and for many helpful discussions. This work is supported by the Japan Society for the Promotion of Science (JSPS) through the Research Fellowship for Young Scientists (DC1), Grant Number JP25KJ1664.
2 Well known basics
In this section, we summarize basic facts about $\mathfrak{gl}(m|n)$ and its representation theory. All material in this section is well known.
Let the base field $k$ be an algebraically closed field of characteristic 0. Let $\mathbf{Vec}$ denote the category of vector spaces, and $\mathbf{sVec}$ the category of supervector spaces. Let $\Pi:\mathbf{sVec}\to\mathbf{sVec}$ be the parity–shift functor. Let $F:\mathbf{sVec}\to\mathbf{Vec}$ be the monoidal functor that forgets the $\mathbb{Z}/2\mathbb{Z}$–grading. Note that $F$ is not a symmetric monoidal functor. In what follows, whenever we refer to the dimension of a supervector space, we mean this total (ordinary) dimension:
$\dim V:=\dim_{k}F(V).$
From now on, we will denote by $\mathfrak{g}$ a finite dimensional Lie superalgebra. We consider the category $\mathfrak{g}\text{-sMod}$, where morphisms respects $\mathbb{Z}/2\mathbb{Z}$-grading. (This is the module category of a monoid object in $\mathfrak{g}\text{-sMod}$ in the sense of [etingof2015tensor].)
2.1 Root systems $\Delta$
Let $V=V_{\overline{0}}\oplus V_{\overline{1}}$ be a $\mathbb{Z}/2\mathbb{Z}$-graded vector space, where $V_{\overline{0}}$ (the even part) is spanned by $v_{1},\dots,v_{m}$ and $V_{\overline{1}}$ (the odd part) is spanned by $v_{m+1},\dots,v_{m+n}$.
The space $\operatorname{End}(V)$ is spanned by basis elements $E_{ij}$, defined by: $E_{ij}\cdot v_{k}=\delta_{jk}v_{i}.$
The general linear Lie superalgebra $\mathfrak{gl}(m|n)$ is defined as the Lie superalgebra spanned by all $E_{ij}$ with $1\leq i,j\leq m+n$, under the supercommutator: $E_{ij},E_{kl}]=E_{ij}E_{kl}-(-1)^{|E_{ij}||E_{kl}|}E_{kl}E_{ij},$) where $|E_{ij}|=\overline{0}$ if $E_{ij}$ acts within $V_{\overline{0}}$ or $V_{\overline{1}}$ (even), and $|E_{ij}|=\overline{1}$ if it maps between $V_{\overline{0}}$ and $V_{\overline{1}}$ (odd).
In the general linear Lie superalgebra $\mathfrak{g}=\mathfrak{gl}(m|n)$, the even part is given by $\mathfrak{g}_{\overline{0}}=\mathfrak{gl}(m)\oplus\mathfrak{gl}(n).$
We fix the standard Cartan subalgebra $\mathfrak{h}:=\bigoplus_{1\leq i\leq m+n}kE_{ii}.$ Define linear functionals $\varepsilon_{1},\dots,\varepsilon_{n+m}\in\mathfrak{h}^{*}$ by requiring that $\varepsilon_{i}(E_{jj})=\delta_{ij}$ for $1\leq i,j\leq n+m$. Define $\delta_{i}=\varepsilon_{m+i}$ for $1\leq i\leq n$. The non-degenerate symmetric bilinear form (,)(\,,\,) on is defined as follows: $(\varepsilon_{i},\varepsilon_{j})=(-1)^{|v_{i}|}\delta_{i,j}$
The root space $\mathfrak{g}_{\alpha}$ associated with $\alpha\in\mathfrak{h}^{*}$ is defined as $\mathfrak{g}_{\alpha}:=\{\,x\in\mathfrak{g}\mid[h,x]=\alpha(h)x,\text{ for all }h\in\mathfrak{h}\,\}.$ Then we have $\mathfrak{g}_{\varepsilon_{i}-\varepsilon_{j}}=kE_{ij}.$
The set of roots $\Delta$ is defined as $\Delta:=\{\,\alpha\in\mathfrak{h}^{*}\mid\mathfrak{g}_{\alpha}\neq 0\,\}\setminus\{0\}.$
We have a root space decomposition of $\mathfrak{g}$ with respect to $\mathfrak{h}$:
$\mathfrak{g}=\mathfrak{h}\oplus\bigoplus_{\alpha\in\Delta}\mathfrak{g}_{\alpha},\qquad\text{and}\quad\mathfrak{g}_{0}=\mathfrak{h}.$
we define The sets of all even roots and odd roots as follows: $\Delta_{\overline{0}}=\{\,\varepsilon_{i}-\varepsilon_{j},\ \delta_{i}-\delta_{j}\mid i\neq j\,\},$ $\Delta_{\overline{1}}=\{\,\varepsilon_{i}-\delta_{j}\mid 1\leq i\leq m,\ 1\leq j\leq n\,\}.$ Note that for any odd root $\alpha\in\Delta_{\bar{1}}$ we have $(\alpha,\alpha)=0$; i.e., every odd root is isotropic .
Definition 2.1.
A subset $\Delta^{+}\subset\Delta$ is called a positive system if it satisfies:
  1. 1.
    $\Delta=\Delta^{+}\cup(-\Delta^{+})$ (disjoint union),
  2. 2.
    For any $\alpha,\beta\in\Delta^{+}$, if $\alpha+\beta\in\Delta$, then $\alpha+\beta\in\Delta^{+}$.
Given a positive system $\Delta^{+}$, the associated fundamental system is defined by $\Pi:=\{\alpha\in\Delta^{+}\mid\alpha\text{ cannot be written as }\alpha=\beta+\gamma\text{ with }\beta,\gamma\in\Delta^{+}\}.$ Elements of $\Pi$ are called simple roots.
Let $\Delta^{+}\subset\Delta$ be a positive system. The corresponding Borel subalgebra $\mathfrak{b}\subset\mathfrak{g}$ is defined as $\mathfrak{b}:=\mathfrak{h}\oplus\bigoplus_{\alpha\in\Delta^{+}}\mathfrak{g}_{\alpha},$ where $\mathfrak{g}_{\alpha}\subset\mathfrak{g}$ is the root space corresponding to $\alpha\in\Delta$. Choose natural root vectors $e_{\alpha}\in\mathfrak{g}_{\alpha}$, $e_{-\alpha}\in\mathfrak{g}_{-\alpha}$ Then for any $\lambda\in\mathfrak{h}^{*}$, $\lambda([e_{\alpha},e_{-\alpha}])=(\lambda,\alpha).$
For a Borel subalgebra $\mathfrak{b}$, we express the triangular decomposition of $\mathfrak{g}$ as $\mathfrak{g}=\mathfrak{n}^{\mathfrak{b}-}\oplus\mathfrak{h}\oplus\mathfrak{n}^{\mathfrak{b}+},$ where $\mathfrak{b}=\mathfrak{h}\oplus\mathfrak{n}^{\mathfrak{b}+}$.
The sets of positive roots, even positive roots, and odd positive roots corresponding to $\mathfrak{b}$ are denoted by $\Delta^{\mathfrak{b},+}$, $\Delta_{\bar{0}}^{\mathfrak{b},+}$, and $\Delta_{\bar{1}}^{\mathfrak{b},+}$, respectively. Let $\Pi^{\mathfrak{b}}$ denote the set of all simple roots associated with the fixed Borel subalgebra $\mathfrak{b}$, and write $\Pi^{\mathfrak{b}}=\Pi^{\mathfrak{b}}_{\bar{0}}\sqcup\Pi^{\mathfrak{b}}_{\bar{1}}$ for its decomposition into even and odd simple roots.
It is clear that the root system of $\mathfrak{gl}(m|n)$, after forgetting the parity of roots, identifies with that of $\mathfrak{gl}(m+n)$. In particular, it is useful to consider the Weyl group $S_{m+n}$ of $\mathfrak{gl}(m+n)$. This phenomenon is specific to $\mathfrak{gl}(m|n)$ and does not occur for other basic classical Lie superalgebras.
The Weyl group $W:=S_{m}\times S_{n}$ of $\mathfrak{g}_{\bar{0}}=\mathfrak{gl}(m)\oplus\mathfrak{gl}(n)$ acts on $\mathfrak{h}^{*}$ by permuting the $\varepsilon_{i}$’s and the $\delta_{j}$’s separately. For an even root $\alpha$, we denote by $s_{\alpha}$ the corresponding reflection in $W$. We denote by $w_{0}$ the longest element of the Weyl group $W$.
The even positive root systems are all conjugate under $S_{m}\times S_{n}$. We therefore fix the standard even Borel $\mathfrak{b}^{\mathrm{st}}_{\bar{0}}$ (block upper triangular in $\mathfrak{g}_{\bar{0}}$), with even positive root system
$\Delta_{\bar{0}}^{+}=\{\ \varepsilon_{i}-\varepsilon_{j}\mid 1\leq i<j\leq m\ \}\ \cup\ \{\ \delta_{p}-\delta_{q}\mid 1\leq p<q\leq n\ \}.$
With this choice fixed, positive systems whose even part equals $\Delta_{\bar{0}}^{+}$ are classified by $\varepsilon\delta$–sequences, equivalently by $(m|n)$-shuffle
$\tau\in\mathrm{Sh}(m|n):=\bigl\{\,w\in S_{m+n}\ \big|\ w(1)<\cdots<w(m)\ \text{and}\ w(m+1)<\cdots<w(m+n)\,\bigr\}.$
Given $\tau$, set
$\Delta^{+}(\tau)=\{\ \varepsilon_{i}-\varepsilon_{j}\mid\tau(i)<\tau(j)\ \}$
yielding a bijection $\tau\leftrightarrow\Delta^{+}(\tau)$; hence there are $\frac{(m+n)!}{m!n!}$ positive systems with standard even one.
Definition 2.2.
A partition is a weakly decreasing finite sequence of positive integers; we allow the empty partition (). For example, we write $5+3+3+1$ as $(53^{2}1)$. We identify partitions with Young diagrams in French notation (rows increase downward). A box has coordinates $(i,j)$ with $i$ the column (from the left) and $j$ the row (from the bottom).
Each $\varepsilon\delta$–sequences with $n$ symbols $\varepsilon$ and $m$ symbols $\delta$ determines a lattice path from $(n,0)$ to $(0,m)$ (left step for $\varepsilon$, up step for $\delta$); the region weakly southeast of the path inside $(n^{m})$ is a Young diagram fitting in the $m\times n$ rectangle $(n^{m})$, and this gives a bijection between $\varepsilon\delta$–sequences and the set of Young diagrams fitting in the $m\times n$ rectangle $(n^{m})$.
In what follows, we represent a Borel subalgebra with standard even Borel subalgebra by a partition. In particular, we call () the distinguished Borel subalgebra, and $(n^{m})$ the anti-distinguished Borel subalgebra. We denote by $L(m,n)$ the set of all Borel subalgebras of $\mathfrak{gl}(m|n)$ whose even part is the standard Borel subalgebra of $\mathfrak{g}_{\overline{0}}$.
2.2 Contragredient duality $()^{\vee}$
Definition 2.3.
A weight $\lambda\in\mathfrak{h}^{*}$ is integral if $\lambda=\sum_{i=1}^{m}a_{i}\,\varepsilon_{i}+\sum_{j=1}^{n}b_{j}\,\delta_{j}$ with $a_{i},b_{j}\in\mathbb{Z}$ for all $i,j$. Write the set of integral weights as $\Lambda$.
We denote by $s\mathcal{W}$ the full subcategory of $\mathfrak{g}$-modules which are $\mathfrak{h}$-semisimple with integral weights and finite-dimensional weight spaces (i.e. integral weight modules with finite-dimensional weight spaces).
A module $M\in s\mathcal{W}$ admits a weight space decomposition :
$M=\bigoplus_{\lambda\in\Lambda}M_{\lambda},\qquad M_{\lambda}:=\{v\in M\mid h\cdot v=\lambda(h)v\ \text{for all }h\in\mathfrak{h}\}.$
Lemma 2.4 (See also [brundan2014representations] Lemma 2.2, [chen2020primitive] Proposition 2.2.3).
We can choose $p\in\operatorname{Map}(\Lambda,\mathbb{Z}/2\mathbb{Z})$ such that
$\mathcal{W}:=\{M\in s\mathcal{W}\mid\deg M_{\lambda}=p(\lambda)\text{ for }\lambda\in\Lambda,M_{\lambda}\neq 0\}$
forms a Serre subcategory, $\mathcal{W}$ contains the trivial module and $s\mathcal{W}=\mathcal{W}\oplus\Pi\mathcal{W}$.
Similarly, for $\mathfrak{g}_{\bar{0}}$ we introduce the category $s\mathcal{W}_{\bar{0}}$, and define the category $\mathcal{W}_{\bar{0}}$ so as to be compatible with the restriction functor $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}$.
Lemma 2.5 ([bell1993theory], Th. 2.2; [coulembier2017gorenstein], §6.1).
Let $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}:\mathfrak{g}\text{-sMod}\to\mathfrak{g}_{\bar{0}}\text{-sMod}$, $\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$, and $\operatorname{Coind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$ denote restriction, induction, and coinduction, respectively; thus $\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\dashv\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\dashv\operatorname{Coind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}$.
Then: (1) each of $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}$, $\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}$, $\operatorname{Coind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}$ is exact and restricts between $\mathcal{W}_{\bar{0}}$ and $\mathcal{W}$; (2) there is a natural isomorphism of functors $\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\cong\operatorname{Coind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}$; (3) The unit $\operatorname{Id}_{\mathcal{O}_{\bar{0}}}\!\to\!\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\!\circ\!\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}$ is a split monomorphism, and the counit $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\!\circ\!\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\twoheadrightarrow\operatorname{Id}_{\mathcal{W}_{\bar{0}}}$ is a split epimorphism. (4) $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\circ\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\cong\Lambda(\mathfrak{g}_{\bar{1}})\otimes$- on $\mathcal{W}_{\bar{0}}$.
Note that $\Lambda^{\dim\mathfrak{g}_{\overline{1}}}(\mathfrak{g}_{\overline{1}})\cong\Lambda^{2mn}(\mathfrak{g}_{\overline{1}})\cong L_{\bar{0}}(0)$ in our general linear setting, but in general, we have $\operatorname{Ind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\cong\operatorname{Coind}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\circ(-\otimes\Lambda^{\dim\mathfrak{g}_{\bar{1}}}\mathfrak{g}_{\bar{1}})$. $L_{\bar{0}}(0)=\Lambda^{0}(\mathfrak{g}_{\overline{1}})$ and $\Lambda^{\dim\mathfrak{g}_{\overline{1}}}(\mathfrak{g}_{\overline{1}})$ are direct summands of $\Lambda(\mathfrak{g}_{\overline{1}})$.
Definition 2.6.
For a module $M$ in the category $\mathcal{W}$, the character $\operatorname{ch}M$ is a formal sum that encodes the dimensions of the weight spaces of $M$: $\operatorname{ch}M:=\sum_{\lambda}\dim M_{\lambda}e^{\lambda}$.
Lie superalgebra $\mathfrak{gl}(m|n)$ has an antiautomorphism $\tau$ given by the formula $\tau(E_{ij}):=-(-1)^{|i|(|i|+|j|)}\,E_{ji}$. For a weight module $M=\bigoplus_{\nu\in\mathfrak{h}^{*}}M_{\nu}$, we let $M^{\vee}:=\bigoplus_{\nu}M_{\nu}^{*}$ be the restricted dual of $M$. We may define an action of $\mathfrak{gl}(m|n)$ on $M^{\vee}$ by $(g\cdot f)(x):=-\,f(\tau(g)x)$, $f\in M^{\vee}$, $g\in\mathfrak{gl}(m|n)$, $x\in M$. This defines a contravariant duality functor $(-)^{\vee}:\mathcal{W}\to\mathcal{W}$.
Proposition 2.7 (Properties of contragredient duality).
  1. 1.
    $(-)^{\vee}$ is an exact contravariant functor.
  2. 2.
    There is a natural isomorphism $M\xrightarrow{\sim}M^{\vee\vee}$ for all $M\in\mathcal{W}$ .
  3. 3.
    $\operatorname{ch}(M^{\vee})=\operatorname{ch}(M)$; in particular, the weight multiplicities are preserved.
  4. 4.
    Since $\tau$ is an automorphism of $\mathfrak{g}$ compatible with the inclusions of subalgebras, there is a natural isomorphism of functors $\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\cong(\,\cdot\,)^{\vee}\circ\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\circ(\,\cdot\,)^{\vee}$. Consequently, $\mathrm{Ind}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\cong(\,\cdot\,)^{\vee}\circ\mathrm{Ind}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\circ(\,\cdot\,)^{\vee}$.
2.3 $\mathfrak{b}$-Verma modules $M^{\mathfrak{b}}(\lambda)$
Let $\mathfrak{b}\in L(m,n)$. Let $k_{\lambda}^{\mathfrak{b}}$ be the one-dimensional $\mathfrak{b}$-module corresponding to $\lambda\in\Lambda$. The $\mathfrak{b}$-Verma module with highest weight $\lambda$ is defined by $M^{\mathfrak{b}}(\lambda)=U(\mathfrak{g})\otimes_{U(\mathfrak{b})}k_{\lambda}^{\mathfrak{b}}.$ (Here, the parity of $k_{\lambda}^{\mathfrak{b}}$ is chosen so that $M^{\mathfrak{b}}(\lambda)\in\mathcal{W}$. )
For the $\mathfrak{b}$-Verma supermodule $M^{\mathfrak{b}}(\lambda)$, we denote its $\mathfrak{b}$-highest weight vector by $v_{\lambda}^{\mathfrak{b}}$. Its simple top is denoted by $L^{\mathfrak{b}}(\lambda)$. Similarly, for the even part $\mathfrak{g}_{\overline{0}}$, the corresponding Verma module and simple module are denoted by $M_{\overline{0}}(\lambda)$ and $L_{\overline{0}}(\lambda)$, respectively.
We define Berezinian weight $ber:=\varepsilon_{1}+\cdots+\varepsilon_{m}-(\delta_{1}+\cdots+\delta_{n}),$ A weight $\lambda$ is orthogonal to all roots if and only if $\lambda=tber$ for some $t\in k$ (i.e. a scalar multiple of the Berezinian weight.
Definition 2.8 (integral Weyl vectors [brundan2014representations]).
Define the vectors
$\rho_{\bar{0}}:=\frac{1}{2}\sum_{\beta\in\Delta_{\bar{0}}^{+}}\beta,\qquad\rho_{\bar{1}}^{\mathfrak{b}}:=\frac{1}{2}\sum_{\gamma\in\Delta_{\bar{1}}^{\mathfrak{b},+}}\gamma,$
and write $\rho_{\bar{1}}:=\rho_{\bar{1}}^{()}$. We have $\rho_{\bar{1}}=-\rho_{\bar{1}}^{(n^{m})}$.
Let the integral Weyl vector be
$\rho:=-\varepsilon_{2}-2\varepsilon_{3}-\cdots-(m-1)\varepsilon_{n}+(m-1)\delta_{m+1}+(m-2)\delta_{m+2}+\cdots+(m-n)\delta_{m+n}.$
One checks that $\rho=\rho_{\bar{0}}-\rho_{\bar{1}}^{()}-\frac{1}{2}(m+n-1)\,\mathrm{ber}.$
For a Borel $\mathfrak{b}$, set
$\rho^{\mathfrak{b}}:=\rho_{\bar{0}}-\rho_{\bar{1}}^{\mathfrak{b}}-\frac{1}{2}(m+n-1)\,\mathrm{ber}.$
It is worth noting that for a $\mathfrak{b}$-simple root $\alpha$, $\rho^{\,r_{\alpha}\mathfrak{b}}=\rho^{\mathfrak{b}}+\alpha.$ Moreover, if $\alpha\in\Pi^{\mathfrak{b}}$, then $(\rho^{\mathfrak{b}},\alpha)=\frac{1}{2}(\alpha,\alpha).$ In particular, if $\alpha$ is odd isotropic, then $(\alpha,\alpha)=0$ and hence $(\rho^{\mathfrak{b}},\alpha)=0.$
It is convenient to encode an integral weight $\lambda$ by the $m\mid n$-tuple $(\lambda_{1},\dots,\lambda_{m}\mid\lambda_{m+1},\dots,\lambda_{m+n})$ of integers defined by $\lambda_{i}:=(\lambda+\rho,\varepsilon_{i})$. Note that our $\rho^{\mathfrak{b}}$ is integral.
We call $\lambda=(\lambda_{1},\dots,\lambda_{m}\mid\lambda_{m+1},\dots,\lambda_{m+n})$ (or $L^{()}(\lambda)$) regular iff $\lambda_{i}\neq\lambda_{j}$ for all $1\leq i<j\leq m$ and $m+1\leq i<j\leq m+n$.
For an integral weight $\lambda=(\lambda_{1},\dots,\lambda_{m}\mid\lambda_{m+1},\dots,\lambda_{m+n})$, $\lambda$ (or $L^{()}(\lambda)$) is antidominant iff $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{m}$ and $\lambda_{m+1}\geq\lambda_{m+2}\geq\cdots\geq\lambda_{m+n}$.
We say that $\lambda$ (or $L^{()}(\lambda)$) is dominant iff $\lambda_{1}\geq\cdots\geq\lambda_{m}$ and $\lambda_{m+1}\leq\cdots\leq\lambda_{m+n}$ .
For $w\in W$ and a weight $\lambda\in\Lambda$, we define the $\mathfrak{b}$-dot action of $W$ on $\Lambda$ by
$w\cdot_{\mathfrak{b}}\lambda:=w(\lambda+\rho^{\mathfrak{b}})-\rho^{\mathfrak{b}}.$
When $\mathfrak{b}=()$, we simply write $w\cdot\lambda$ for $w\cdot_{()}\lambda$.
The (usual) dot action of $W$ on $\Lambda$ is defined by
$w\cdot\lambda:=w(\lambda+\rho_{\overline{0}})-\rho_{\overline{0}}.$
$\rho^{\mathfrak{b}}_{\overline{1}}$ is WW–invariant iff $\mathfrak{b}=(),(n^{m})$ , so the three actions coincide:
$w\cdot_{()}\lambda=w\cdot_{(n^{m})}\lambda=w\cdot\lambda\qquad\text{for all }w\in W,\ \lambda\in\Lambda.$
For $\lambda\in\Lambda$, we denote by $\lambda^{\mathrm{dom}}$ and $\lambda^{\mathrm{antidom}}$ the unique dominant and antidominant weights, respectively, in the dot-orbit $W\cdot\lambda$.
We define the $\mathfrak{b}$-atypicality of $\lambda$ as
[…]
If $\mathrm{aty}^{\mathfrak{b}}(\lambda)=0$, then $\lambda$ is called $\mathfrak{b}$-typical; otherwise, it is called $\mathfrak{b}$-atypical.
[…]
For a root $\alpha$, we write $E_{\alpha}$ for the natural root vector corresponding to $\alpha$.
Lemma 2.9 ([musson2012lie]).
Let $\mathfrak{g}=\mathfrak{gl}(m|n)$ and let $\mathfrak{b}$ be a Borel subalgebra. Let $\alpha$ be an even $\mathfrak{b}$-simple root . For an integral weight $\lambda\in\Lambda$ , if $\langle\lambda+\rho^{\mathfrak{b}},\alpha^{\vee}\rangle=m\in\mathbb{Z}_{>0}$ , then $\dim\operatorname{Hom}\bigl(M^{\mathfrak{b}}(s_{\alpha}\cdot_{\mathfrak{b}}\lambda),M^{\mathfrak{b}}(\lambda)\bigr)=1$ , and any nonzero homomorphism is injective. It is given by $v_{s_{\alpha}\cdot_{\mathfrak{b}}\lambda}^{\mathfrak{b}}\mapsto{E_{-\alpha}^{m}}v_{\lambda}^{\mathfrak{b}}$
Theorem 2.10 ([musson2012lie]).
Let $\mathfrak{b}$ be a Borel subalgebra. Let $\lambda,\mu\in\Lambda$ be weights such that $\lambda-\mu\;\in\;\sum_{\alpha\in\Delta_{\overline{0}}^{+}}\mathbb{Z}_{\geq 0}\,\alpha.$ Then
$\dim\operatorname{Hom}_{\mathfrak{g}}\bigl(M^{\mathfrak{b}}(\mu),\,M^{\mathfrak{b}}(\lambda)\bigr)\;\leq\;1,$
and every nonzero homomorphism $M^{\mathfrak{b}}(\mu)\longrightarrow M^{\mathfrak{b}}(\lambda)$ is injective.
Lemma 2.11.
For $\mathfrak{b},\mathfrak{b}^{\prime}\in L(m,n)$ and $\lambda,\lambda^{\prime}\in\Lambda$, the following statements hold:
  1. 1.
    $\operatorname{ch}M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}})=\operatorname{ch}M^{\mathfrak{b}^{\prime}}(\lambda-\rho^{\mathfrak{b}^{\prime}})$;
  2. 2.
    $\dim M^{\mathfrak{b}^{\prime}}(\lambda-\rho^{\mathfrak{b}^{\prime}})_{\lambda-\rho^{\mathfrak{b}}}=1$;
  3. 3.
    $\operatorname{ch}M^{\mathfrak{b}}(\lambda)=\operatorname{ch}M^{\mathfrak{b}^{\prime}}(\lambda^{\prime})\iff\lambda+\rho^{\mathfrak{b}}=\lambda^{\prime}+\rho^{\mathfrak{b}^{\prime}}$;
  4. 4.
    $\dim\operatorname{Hom}(M^{\mathfrak{b}}(\lambda-\rho^{\mathfrak{b}}),M^{\mathfrak{b}^{\prime}}(\lambda-\rho^{\mathfrak{b}^{\prime}}))=1$;
2.4 Kac functors $K_{\geq 0}$
Recall that we represent a Borel subalgebrawith sandard even Borel subalgebra by a partition. In particular, we denote () the distinguished Borel subalgebra, and $(n^{m})$ the anti-distinguished Borel subalgebra.
Define $\mathfrak{g}_{1}:=()_{\bar{1}}\qquad\mathfrak{g}_{-1}:=(n^{m})_{\bar{1}}.$
The general linear Lie superalgebra $\mathfrak{gl}(m|n)$ admits a natural $\mathbb{Z}$-grading given by $\mathfrak{gl}(m|n)=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{1}.$ This grading satisfies the relations $\mathfrak{g}_{1},\mathfrak{g}_{1}]=[\mathfrak{g}_{-1},\mathfrak{g}_{-1}]=0.$ We define $\mathfrak{g}_{\geq 0}:=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{1},\quad\mathfrak{g}_{\leq 0}:=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{\bar{0}}.$
Define the inflation functor $\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}\colon\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}_{\geq 0}\text{-sMod}$ by letting $\mathfrak{g}_{1}$ act trivially. Define the functor $(\,)^{\mathfrak{g}_{1}}\colon\mathfrak{g}_{\geq 0}\text{-sMod}\to\mathfrak{g}_{\bar{0}}\text{-sMod}$ by $M\mapsto M^{\mathfrak{g}_{1}}:=\{\,m\in M\mid\mathfrak{g}_{1}m=0\,\}$ . Then these functors are mutually adjoint each other and exact.
The Kac functor is defined as $K_{\geq 0}:=\mathrm{Ind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}\circ\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$, which is exact. Then $K_{\geq 0}$ is left adjoint to the exact functor $\mathrm{Res}_{\geq 0}:=(\,)^{\mathfrak{g}_{1}}\circ\mathrm{Res}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}:\mathfrak{g}\text{-sMod}\to\mathfrak{g}_{\bar{0}}\text{-sMod}$. If we denote by $\mathrm{Coind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}$ [chen2021simple] the right adjoint of $\mathrm{Res}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}$, and define $\mathrm{Coind}_{\geq 0}:=\mathrm{Coind}_{\mathfrak{g}_{\geq 0}}^{\mathfrak{g}}\circ\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\geq 0}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$ then there is an isomorphism of functors $\mathrm{Coind}_{\geq 0}\circ(-\otimes L_{\bar{0}}(-2\rho_{\bar{1}}))\cong K_{\geq 0}$, see [chen2021simple]. Note that $L_{\bar{0}}(-2\rho_{\bar{1}})\cong\bigwedge^{\dim\mathfrak{g}_{-1}}\mathfrak{g}_{-1}$ as $\mathfrak{g}_{\bar{0}}$-modules.
Similarly, we define the dual Kac functor $K_{\leq 0}:=\mathrm{Ind}_{\mathfrak{g}_{\leq 0}}^{\mathfrak{g}}\circ\mathrm{Infl}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}_{\leq 0}}:\mathfrak{g}_{\bar{0}}\text{-sMod}\to\mathfrak{g}\text{-sMod}$, $\mathrm{Res}_{\leq 0}:=\mathrm{Res}_{\mathfrak{g}_{\leq 0}}^{\mathfrak{g}}\circ(\,\cdot\,)^{\mathfrak{g}_{-1}}$ and $\mathrm{Coind}_{\leq 0}$.
Lemma 2.12.
[germoni1998indecomposable] $K_{\geq 0}\cong(\,\cdot\,)^{\vee}\circ\mathrm{Coind}_{\leq 0}\circ(\,\cdot\,)^{\vee},\qquad K_{\leq 0}\cong(\,\cdot\,)^{\vee}\circ\mathrm{Coind}_{\geq 0}\circ(\,\cdot\,)^{\vee}.$ $K_{\geq 0}$ and $K_{\leq 0}$ restricts between $\mathcal{O}_{\bar{0}}$ and $\mathcal{O}$.
Proposition 2.13.
For every weight $\lambda$ one has $K_{\geq 0}\big(M_{\bar{0}}(\lambda)\big)\cong M^{()}(\lambda)$ and $K_{\leq 0}\big(M_{\bar{0}}(\lambda)\big)\cong M^{(n^{m})}(\lambda)$. Using $(\,\cdot\,)^{\vee}$, we have $\mathrm{Coind}_{\leq 0}\big(M_{\bar{0}}(\lambda)^{\vee}\big)\cong K_{\leq 0}\big(M_{\bar{0}}(\lambda-2\rho_{\bar{1}})^{\vee}\big)\cong M^{()}(\lambda)^{\vee}$, and $\mathrm{Coind}_{\geq 0}\big(M_{\bar{0}}(\lambda)^{\vee}\big)\cong K_{\geq 0}\big(M_{\bar{0}}(\lambda+2\rho_{\bar{1}})^{\vee}\big)\cong M^{(n^{m})}(\lambda)^{\vee}$.
Proposition 2.14.
( [chen2021simple]) For a weight $\lambda$, the following are true:
  1. 1.
    $K_{\geq 0}(L_{\bar{0}}(\lambda))^{\vee}\cong K_{\leq 0}(L_{\bar{0}}(\lambda-2\rho_{\bar{1}}))$,
  2. 2.
    $\operatorname{soc}K_{\geq 0}(L_{\bar{0}}(\lambda))\cong L^{(n^{m})}(\lambda-2\rho_{\bar{1}})$,
  3. 3.
    $\operatorname{top}K_{\geq 0}(L_{\bar{0}}(\lambda))\cong L^{()}(\lambda)$;
  4. 4.
    $\dim\operatorname{Hom}(K_{\geq 0}(L_{\bar{0}}(\lambda)),K_{\leq 0}(L_{\bar{0}}(\mu)))=\delta_{\lambda,\mu+2\rho_{\bar{1}}}$;
  5. 5.
    $K_{\geq 0}(L_{\bar{0}}(\lambda))\cong K_{\leq 0}(L_{\bar{0}}(\lambda-2\rho_{\bar{1}})))$ iff $K_{\geq 0}(L_{\bar{0}}(\lambda))\cong L(\lambda)$ iff $L(\lambda)$ is typical.
Theorem 2.15.
[chen2021translated] For every $M\in\mathfrak{g}_{\bar{0}}\text{-sMod}$ which is finite length, one has $\ell(\operatorname{soc}M)=\ell(\operatorname{soc}K_{\geq 0}M)=\ell(\operatorname{soc}K_{\leq 0}M)$ and $\ell(\operatorname{top}M)=\ell(\operatorname{top}K_{\geq 0}M)=\ell(\operatorname{top}K_{\leq 0}M)$, where $\ell(-)$ denotes the length.
In particular, for every weight $\lambda\in\Lambda$, $\operatorname{soc}M^{()}(\lambda)$ is simple.
Proposition 2.16 ([mazorchuk2014parabolic, chen2021tilting]).
For a weight $\lambda$, the following are equivalent:
0. $\lambda$ is antidominant.
  1. 1.
    $M_{\bar{0}}(\lambda)\cong L_{\bar{0}}(\lambda)\cong M_{\bar{0}}(\lambda)^{\vee}$.
  2. 2.
    $P_{\bar{0}}(\lambda)\cong I_{\bar{0}}(\lambda)$.
  3. 3.
    $M^{()}(\lambda)\cong K_{\geq 0}(L_{\bar{0}}(\lambda))$.
  4. 4.
    $M^{()}(\lambda)\cong M^{(n^{m})}(\lambda-2\rho_{\bar{1}})^{\vee}$.
Theorem 2.17 ([chen2023some]).
labelthm:soc-dist Let $\lambda$ be an integral weight . Then
$\operatorname{Soc}M^{()}(\lambda)\;\cong\;L^{(n^{m})}\bigl(\lambda^{\mathrm{antidom}}-2\rho_{\overline{1}}\bigr).$
Proof. There exists an embedding
$M^{()}(\lambda^{\mathrm{antidom}})\hookrightarrow M^{()}(\lambda).$
Hence
$\operatorname{Soc}M^{()}(\lambda)\supset\operatorname{Soc}M^{()}(\lambda^{\mathrm{antidom}})\;\cong\;L^{(n^{m})}\bigl(\lambda-2\rho_{\overline{1}}\bigr),$
where the last isomorphism follows from  2.4. On the other hand, by Theorem 2.15 , the socle $\operatorname{Soc}M^{()}(\lambda)$ is simple. $\square$
Corollary 2.18.
Let $\lambda,\mu\in\Lambda$ be integral weights. If there exists an embedding
$M^{()}(\mu)\;\hookrightarrow\;M^{()}(\lambda),$
then $\mu\in W\cdot\lambda$.
2.5 $\mathfrak{g}_{-1}$-generic weights
To each regular dominant weight $\lambda\in\Lambda$ (or $L^{()}(\lambda)$) we associate, following [stroppel2012highest], its weight diagram as following.
$I_{\times}(\lambda):=\{\lambda_{1},\lambda_{2},\ldots,\lambda_{m}\}$ and $I_{\circ}(\lambda):=\{\lambda_{m+1},\lambda_{m+2},\ldots,\lambda_{m+n}\}.$ The integers in $I_{\times}(\lambda)\cap I_{\circ}(\lambda)$ are labelled by $\vee$, the remaining ones in $I_{\times}(\lambda)$ respectively $I_{\circ}(\lambda)$ are labelled by $\times$ respectively $\circ$. All other integers are labelled by a $\wedge$. This labelling of the number line $\mathbb{Z}$ uniquely characterizes the weight $\lambda$.
Definition 2.19 (Totally disconnected weights).
A regular dominant weight $\lambda\in\Lambda$ (or, equivalently, the simple module $L^{()}(\lambda)$) is called totally disconnected if its weight diagram contains at least one $\wedge$” between any two $\vee$”s.
Definition 2.20 ($\mathfrak{g}_{-1}$-generic weights).
A weight $\lambda\in\Lambda$ (or $L^{()}(\lambda)$) is called $\mathfrak{g}_{-1}$-generic if the set
$\bigl\{\,\lambda-\textstyle\sum_{\alpha\in S}\alpha\,\bigm|\,S\subset\Delta_{\overline{1}}^{()+}\,\bigr\}$
lie in the same $\rho$–shifted Weyl chamber.
Lemma 2.21.
Let $\lambda\in\Lambda$ be $\mathfrak{g}_{-1}$-generic. Then:
  1. 1.
    $\lambda$ is regular
  2. 2.
    If $\lambda$ is dominant, then $\lambda$ is totally disconnected.
  3. 3.
    For any Borel subalgebra $\mathfrak{b}\in L(m|n)$ we have $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\overline{0}}}M^{\mathfrak{b}}\bigl(\lambda+\rho^{()}-\rho^{\mathfrak{b}}\bigr)\;\cong\;\bigoplus_{S\subset\Delta_{\overline{1}}^{()+}}M_{\overline{0}}\Bigl(\lambda-\sum_{\alpha\in S}\alpha\Bigr).$
  4. 4.
    $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\overline{0}}}K_{\geq 0}\bigl(L_{\overline{0}}(\lambda)\bigr)\;\cong\;\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\overline{0}}}K_{\leq 0}(L_{\bar{0}}(\lambda-2\rho_{\bar{1}}))\;\cong\;\bigoplus_{S\subset\Delta_{\overline{1}}^{+}}L_{\overline{0}}\Bigl(\lambda-\sum_{\alpha\in S}\alpha\Bigr).$
  5. 5.
    For every $\mu\in W\cdot\lambda$ , $\mu$ is again $\mathfrak{g}_{-1}$-generic.
3 Narrow Verma modules $N^{()}(\lambda)$
Definition 3.1 (Narrow Verma modules).
We define the narrow Verma module associated $\lambda$ by
$N^{()}(\lambda)\;:=\;\operatorname{Im}\Bigl(M^{()}(\lambda)\longrightarrow M^{(n^{m})}\bigl(\lambda-2\rho_{\overline{1}}\bigr).\Bigr),$
where we take any nonzero homomorphism $M^{()}(\lambda)\longrightarrow M^{(n^{m})}\bigl(\lambda-2\rho_{\overline{1}}\bigr).$
The following is well known.
Proposition 3.2.
Set $E_{\mathfrak{g}_{1}}:=\prod_{\alpha\in\Delta_{1}^{()+}}E_{\alpha}\in U(\mathfrak{g}_{1})$. Then the following statements hold.
  1. 1.
    The element $E_{\mathfrak{g}_{1}}$ up to scalar, does not depend on the order of the product $\prod_{\alpha\in\Delta_{1}^{()+}}E_{\alpha}$.
  2. 2.
    up to scalar, $x\,E_{\mathfrak{g}_{1}}=E_{\mathfrak{g}_{1}}\,x$ for all $x\in U(\mathfrak{n}_{\overline{0}}^{()-})$.
  3. 3.
    For every $\lambda\in\Lambda$, the narrow Verma module $N^{()}(\lambda)$ is the $\mathfrak{g}$-submodule of $M^{(n^{m})}(\lambda-2\rho_{\overline{1}})$ generated by $E_{\mathfrak{g}_{1}}v^{(n^{m})}_{\lambda-2\rho_{\overline{1}}}$; more precisely,
    $N^{()}(\lambda)=U(\mathfrak{g})\,E_{\mathfrak{g}_{1}}\,v^{(n^{m})}_{\lambda-2\rho_{\overline{1}}}\subset M^{(n^{m})}(\lambda-2\rho_{\overline{1}}).$
Lemma 3.3.
Let $\lambda\in\Lambda$ be an integral weight and let $w\in W$. Assume we have embeddings
$i_{()}\colon M^{()}\bigl(w\cdot\lambda\bigr)\longrightarrow M^{()}(\lambda),$
$i_{(n^{m})}\colon M^{(n^{m})}\bigl(w\cdot\lambda-2\rho_{\overline{1}}\bigr)\longrightarrow M^{(n^{m})}\bigl(\lambda-2\rho_{\overline{1}}\bigr),$
given as compositions of homomorphisms arising from even reflections. Then the following natural diagram commutes:
$M^{()}\bigl(w\cdot\lambda\bigr)$$M^{()}(\lambda)$$M^{(n^{m})}\bigl(\lambda-2\rho_{\overline{1}}\bigr)$$M^{(n^{m})}\bigl(w\cdot\lambda-2\rho_{\overline{1}}\bigr)$$i_{()}$$i_{(n^{m})}$$\neq 0$$\neq 0$
where the horizontal arrows denote the unique (up to scalar) nonzero homomorphisms
$M^{()}\bigl(w\cdot\lambda\bigr)\to M^{(n^{m})}\bigl(w\cdot\lambda-2\rho_{\overline{1}}\bigr),\qquad M^{()}(\lambda)\to M^{(n^{m})}(\lambda-2\rho_{\overline{1}}\bigr).$
Proof. It follows from Section 3 $\square$
Proposition 3.4.
Let $\lambda\in\Lambda$ , then,
$\operatorname{Soc}N^{()}(\lambda)\;\cong\;N^{()}(\lambda^{\mathrm{antidom}})\;\cong\;L^{()}(\lambda^{\mathrm{antidom}}).$
Proposition 3.5.
Let $\lambda\in\Lambda$ be $\mathfrak{g}_{-1}$-generic. Then,
$\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\overline{0}}}N^{()}(\lambda)\quad\text{is a direct summand of}\quad\bigoplus_{S\subset\Delta_{\overline{1}}^{+}}M_{\overline{0}}\Bigl(\lambda-\sum_{\alpha\in S}\alpha\Bigr)$
as a $\mathfrak{g}_{\overline{0}}$-module.
Proof. Apply the restriction functor $\operatorname{Res}_{\overline{0}}$ to a nonzero homomorphism $M^{()}(\lambda)\to M^{(n^{m})}(\lambda-2\rho_{\overline{1}})$. Then by Section 3, $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\overline{0}}}M^{()}(\lambda)$ and $\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\overline{0}}}M^{(n^{m})}(\lambda-2\rho_{\overline{1}})$ decompose as direct sums of $\mathfrak{g}_{\overline{0}}$–Verma modules which are either isomorphic or lie in different blocks. Since the endomorphism ring of a Verma module is one–dimensional and there are no nonzero homomorphisms between Vermas in different blocks, the induced map is just a projection onto (and inclusion of) some direct summands. In particular, on the summands we care about, it is given by taking direct summands. $\square$
4 BGG resolutions
Here, we recalls classical BGG resolution.
Theorem 4.1 (BGG resolution).
Let $\mathfrak{g}=\mathfrak{gl}(m)\oplus\mathfrak{gl}(n)$ and assume that $\lambda$ is regular dominant integral. Then there exists an exact sequence $BGG_{\bar{0}}(\lambda)$ of $\mathfrak{g}$-modules
$0\to M_{\overline{0}}(w_{0}\cdot\lambda)\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=\ell(w_{0})-1\end{subarray}}M_{\overline{0}}(w\cdot\lambda)\to\cdots\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=1\end{subarray}}M_{\overline{0}}(w\cdot\lambda)\to M_{\overline{0}}(\lambda)\to L_{\overline{0}}(\lambda)\to 0.$
The following is our main result.
Theorem 4.2.
Let $\mathfrak{g}=\mathfrak{gl}(m|n)$ and assume that $\lambda$ is dominant integral and $\mathfrak{g}_{-1}$-generic. Then there exists an exact sequence of $\mathfrak{g}$-modules
$0\to N^{()}(w_{0}\cdot\lambda)\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=\ell(w_{0})-1\end{subarray}}N^{()}(w\cdot\lambda)\to\cdots\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=1\end{subarray}}N^{()}(w\cdot\lambda)\to N^{()}(\lambda)\to L^{()}(\lambda)\to 0.$
Proof. Applying the Kac functor to Theorem 4.1, we obtain the following exact sequence of $\mathfrak{g}$-modules $K_{\geq 0}\mathrm{BGG}_{\bar{0}}(\lambda)$:
$0$ $\to M^{()}(w_{0}\cdot\lambda)\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=\ell(w_{0})-1\end{subarray}}M^{()}(w\cdot\lambda)$
$\to\cdots\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=1\end{subarray}}M^{()}(w\cdot\lambda)\to M^{()}(\lambda)\to K_{\geq 0}\bigl(L_{\overline{0}}(\lambda)\bigr)\to 0.$
By Section 3, we obtain the following complex, denoted by $\mathrm{BGG}^{()}(\lambda)$:
$0$ $\to N^{()}(w_{0}\cdot\lambda)\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=\ell(w_{0})-1\end{subarray}}N^{()}(w\cdot\lambda)$
$\to\cdots\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=1\end{subarray}}N^{()}(w\cdot\lambda)\xrightarrow{f}N^{()}(\lambda)\to\operatorname{Cok}(f)\to 0.$
Applying the dual Kac functor to the classical BGG resolution Theorem 4.1, we obtain the following exact sequence of $\mathfrak{g}$-modules $K_{\leq 0}\mathrm{BGG}_{\overline{0}}(\lambda-2\rho_{\overline{1}})$:
$0$ $\to M^{(n^{m})}\bigl(w_{0}\cdot(\lambda-2\rho_{\overline{1}})\bigr)\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=\ell(w_{0})-1\end{subarray}}M^{(n^{m})}\bigl(w\cdot(\lambda-2\rho_{\overline{1}})\bigr)$
$\to\cdots\to\bigoplus_{\begin{subarray}{c}w\in W\\ \ell(w)=1\end{subarray}}M^{(n^{m})}\bigl(w\cdot(\lambda-2\rho_{\overline{1}})\bigr)\to M^{(n^{m})}\bigl(\lambda-2\rho_{\overline{1}}\bigr)\to K_{\leq 0}\bigl(L_{\overline{0}}(\lambda-2\rho_{\overline{1}})\bigr)\to 0.$
By Section 3, we obtain the hom of complex $K_{\geq 0}\mathrm{BGG}_{\bar{0}}(\lambda)\to\mathrm{BGG}^{()}(\lambda)\to K_{\leq 0}\mathrm{BGG}_{\overline{0}}(\lambda-2\rho_{\overline{1}})$. In particular, we obtain following commutative diagram.
$M^{()}(\lambda)$$N^{()}(\lambda)$$M^{(n^{m})}\bigl(\lambda-2\rho_{\overline{1}}\bigr)$$K_{\geq 0}\bigl(L_{\overline{0}}(\lambda)\bigr)$$\operatorname{Cok}(f)$$K_{\leq 0}\bigl(L_{\overline{0}}(\lambda-2\rho_{\overline{1}})\bigr)$000
Applying the restriction functor $\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}$, we obtain a morphism of complexes
$\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\bigl(K_{\geq 0}\mathrm{BGG}_{\overline{0}}(\lambda)\bigr)\;\longrightarrow\;\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\bigl(\mathrm{BGG}^{()}(\lambda)\bigr)\;\longrightarrow\;\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\bigl(K_{\leq 0}\mathrm{BGG}_{\overline{0}}(\lambda-2\rho_{\overline{1}})\bigr).$
By Section 2.5,
$\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\bigl(K_{\geq 0}\mathrm{BGG}_{\overline{0}}(\lambda)\bigr)\;\cong\;\operatorname{Res}^{\mathfrak{g}}_{\mathfrak{g}_{\bar{0}}}\bigl(K_{\leq 0}\mathrm{BGG}_{\overline{0}}(\lambda-2\rho_{\overline{1}})\bigr)\;\cong\;\bigoplus_{S\subset\Delta_{\overline{1}}^{()+}}\mathrm{BGG}_{\overline{0}}\Bigl(\lambda-\sum_{\alpha\in S}\alpha\Bigr).$
By Section 3 and Section 3, the complex $\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\bigl(\mathrm{BGG}^{()}(\lambda)\bigr)$ is a direct summand of both $\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\bigl(K_{\geq 0}\mathrm{BGG}_{\overline{0}}(\lambda)\bigr)$ and $\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\bigl(K_{\leq 0}\mathrm{BGG}_{\overline{0}}(\lambda-2\rho_{\overline{1}})\bigr)$.
Hence the canonical homomorphism
$\operatorname{Cok}(f)\longrightarrow K_{\leq 0}\bigl(L_{\overline{0}}(\lambda-2\rho_{\overline{1}})\bigr)$
is injective.
In particular,
$\operatorname{Cok}(f)\;\cong\;L^{()}(\lambda).$
The exactness of $\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\bigl(\mathrm{BGG}^{()}(\lambda)\bigr)$ follows from the exactness of $\bigoplus_{S\subset\Delta_{\overline{1}}^{()+}}\mathrm{BGG}_{\overline{0}}\Bigl(\lambda-\sum_{\alpha\in S}\alpha\Bigr)$.
Finally, the exactness of $\mathrm{BGG}^{()}(\lambda)$ itself follows from the exactness of $\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}\bigl(\mathrm{BGG}^{()}(\lambda)\bigr)$.
$\square$
Denote by $\Gamma_{\lambda}$ the set of atypical ()-positive roots of $\lambda$:
$\Gamma_{\lambda}=\bigl\{\,\alpha\in\Delta_{\bar{1}}^{()+},\bigm|\,(\lambda+\rho,\ \alpha)=0\bigr\}.$
Theorem 4.3 ([su2007character, chmutov2014weyl]).
Let $\lambda\in\Lambda$ be totally disconnected. Then,
$\operatorname{ch}L^{()}(\lambda)$ $=\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\sum_{w\in W}(-1)^{\ell(w)}\frac{e^{w\cdot\lambda}}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}$
$=\sum_{w\in W}(-1)^{\ell(w)}e^{w\cdot\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}$
$=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
Theorem 4.4.
Let $\lambda\in\Lambda$ be $\mathfrak{g}_{-1}$-generic. Then the character of the narrow Verma module $N^{()}(\lambda)$ is
$\operatorname{ch}N^{()}(\lambda)=e^{\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\\ =\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
Proof. Let $\lambda\in\Lambda$ be dominant. By Section 2.5, the weight $\lambda$ is totally disconnected. Hence, by Theorem 4.3 we have
$\operatorname{ch}L^{()}(\lambda)$ $=\operatorname{ch}\,\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}L^{()}(\lambda)$
$=\sum_{w\in W}(-1)^{\ell(w)}\frac{\operatorname{ch}M^{()}(w\cdot\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}$
$=\sum_{w\in W}(-1)^{\ell(w)}\sum_{S\subset\Delta_{\overline{1}}^{()+}\setminus\Gamma_{\lambda}}\operatorname{ch}M_{\overline{0}}\Bigl(w\cdot\lambda-\sum_{\alpha\in S}\alpha\Bigr).$
On the other hand, by Theorem 4.2 we obtain
$\operatorname{ch}L^{()}(\lambda)=\sum_{w\in W}(-1)^{\ell(w)}\,\operatorname{ch}N^{()}(w\cdot\lambda).$
By Section 3, each $\mathrm{Res}_{\mathfrak{g}_{\bar{0}}}^{\mathfrak{g}}N^{()}(w\cdot\lambda)$ is a direct summand of
$\bigoplus_{S\subset\Delta_{\overline{1}}^{()+}\setminus\Gamma_{w\cdot\lambda}}M_{\overline{0}}\Bigl(w\cdot\lambda-\sum_{\alpha\in S}\alpha\Bigr).$
Note that $w\Gamma_{\lambda}=\Gamma_{w\cdot\lambda}$ for all $w\in W$.
Since the characters $\operatorname{ch}M_{\overline{0}}\bigl(w\cdot\lambda-\sum_{\alpha\in S}\alpha\bigr)$ (for varying $w$) are linearly independent, which yields the desired identity and completes the proof. $\square$
Corollary 4.5.
Let $\lambda\in\Lambda$ be $\mathfrak{g}_{-1}$-generic and antidominant. Then
$\operatorname{ch}L^{()}(\lambda)=e^{\lambda}\,\frac{\displaystyle\prod_{\beta\in\Delta_{\overline{1}}^{()+}\setminus\Gamma_{\lambda}}(1+e^{-\beta})}{\displaystyle\prod_{\gamma\in\Delta_{\overline{0}}^{+}}(1-e^{-\gamma})}\\ =\frac{\displaystyle\operatorname{ch}M^{()}(\lambda)}{\displaystyle\prod_{\beta\in\Gamma_{\lambda}}(1+e^{-\beta})}.$
Shunsuke Hirota
Department of Mathematics, Kyoto University
Kitashirakawa Oiwake-cho, Sakyo-ku, 606-8502, Kyoto
E-mail address: shun299509732@gmail.com